Database last updated on Sun, 01 Feb 2026 21:11:04 +0000 by localhost
Instrument constant changes for the week before 2023298 (2 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 18 rows)
Measurements and tests summary.
| _jday | _loc | _serial | _skc | aode | ds | dt | ed | fm | fz | hg | rs | sl | uv | zs |
| 298 | TORONTO | #014 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.017) | 3 | 3 | 1 | 2 | 32 | 1 | 12 | 1 | 6 | 49 | 63 |
(0 sec)
Number of days since last dome/window cleaning (0 sec, 2 rows)
| serial | days |
| 014 | more than a year |
Missing files (0 sec, 1 rows)
| Serial | last data | days ago |
| 014 | 2023299 | 0-1 |
Questionable Run/Stop results. (0 sec, 1 rows)
| Serial | Jday | slit 0 | dark rs | dark counts | slit 1 | slit 2 | slit 3 | slit 4 | slit 5 | slits 3+5 |
| 014 | 2023298 | 0.999 | | 5 / 176087 (2.5M) | 0.924 | 0.931 | 0.937 | 0.929 | 0.908 | 1.000 |
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 014 | 15 | 2 | 7 | 2 | 2 | 2 | 2 | 2 | 1 |
AP results for MB Brewer. (0 sec, 1 rows)
| Serial | Jday | t1 off | t1 on | t2 off | t2 on | t3 off | t3 on | HV off | HV on | +15V off | +15V on | +5V off | +5V on | -15V off | -15V on | +24V off | +24V on | +5V sps off | +5V sps on | -8V sps off | -8V sps on | SL A off | SL A on | SL V off | SL V on | W |
| 014 | 2023298 | 33.6 | 33.8 | 29.2 | 29.2 | 32.2 | 32.1 | 1549.6 | 1546.5 | 15.1 | 15.0 | 5.1 | 5.1 | -15.1 | -15.1 | 24.1 | 24.1 | 4.9 | 4.9 | -7.9 | -7.9 | 0.1 | 1.6 | 0.0 | 10.93 | 17.3 |
Standard lamp intensity analysis (0 sec, 1 rows)
| Serial | min | max | filter | fixable | 30-day history |
| #014 | 2960580 | 2960580 | 0 | yes |